find the equation to the circle which has its centre at the point (3,4) and touches the straight line,5x+12y-4=0.


Grade:11

1 Answers

ruchi yadav
10 years ago
If straight line touches straight 5x+12y-4=0 so distance of line from center (3,4) will be radius of circle.

so radius will be = (15+48-4)/13 =4.5

so equation will be (x-3)2 + (y-4)2 = (4.5)2


Thank you,
Ruchi
Askiitians Faculty